# Towards Robust Real-World Multivariate Time Series Forecasting: A Unified Framework

### v2

@@ 5.1 Main Results
We evaluate forecasting performance using datasets that reflect realistic sampling heterogeneity across channels. Unlike regular baselines that rely on interpolation, our approach preserves the original sampling structure and operates directly on observed values. Averaging CMSE and CMAE over all prediction lengths (Table 1), our method achieves the best performance on all datasets in the practical setting, highlighting the benefit of avoiding imputation when channels are temporally misaligned. Our method also handles asynchrony more effectively than irregular baselines. Detailed results for each prediction length and dataset are provided in Appendix E.
@@
We further evaluate all models under a more challenging setting, where test-time inputs include block-wise missing intervals, explicitly indicated to the model, on top of channel-wise asynchronous sampling. This configuration encompasses all three core challenges of interest. In this case, all models receive inputs containing contiguous missing segments filled with zeros. But for a fair comparison, linear interpolation is applied to regular baselines to avoid underestimating their capability. Unlike other approaches, our model can internally adjust input length by design, enabling the use of zero-patch masking to effectively handle missing regions. As shown in Table 2, our model consistently outperforms all baselines in average CMSE and CMAE for all prediction lengths across different missing ratios, maintaining high performance even as the severity of missingness increases. This robustness stems from the use of random patch masking during training, which improves the model’s ability to generalize to incomplete patterns. Unlike other methods, our model avoids distortion from artificially zero-filled values and preserves signal fidelity under block-wise missing conditions. Detailed results across datasets are provided in Appendix E. To further demonstrate the overall robustness to missing handling, we conducted additional experiments on the ETT1 and EPA datasets with smaller interval missing settings. The corresponding details are provided in Appendix A.2.
@@ 7 Conclusion
A practical forecasting setting is introduced where three core real-world challenges occur simultaneously: channel-wise asynchronous sampling, block-wise missing intervals, and complex cross-channel dependencies. Instead of interpolation, our approach handles both missing segments and sampling gaps through masking and frequency-based dynamic patching. The unified mask-guided attention leverages channel tokens with read-only masking and cross-channel attention, consistently outperforming prior methods across real-world conditions. Scalability analysis further shows that the masking-based design remains efficient for realistic channel counts and input lengths, although extreme scales may still pose challenges. Beyond scaling, future work will also consider applying ChannelTokenFormer to broader domains, integrating multimodal signals, and improving interpretability of channel interactions, thereby expanding its utility in real-world forecasting.
@@ A.1 Dataset Configuration under Multi-Source Asynchrony
Weather-practical (Weather)
Weather-practical is constructed from the widely used Weather (Wu et al., 2021) benchmark. While the original dataset assumes uniform sampling periods across all channels, this assumption rarely holds in real-world environmental monitoring systems. In Weather-practical, we reorganize the channels to reflect heterogeneous sampling frequencies, based on the intrinsic temporal dynamics of each channel. For instance, fast-changing environmental channels such as wind velocity, precipitation, and solar radiation are sampled at 10-minute intervals, whereas more stable channels such as air pressure or vapor pressure are sampled at a coarser hourly rate. Mid-range channels—such as air temperature and specific humidity—are sampled at 20-minute or 30-minute intervals in accordance with their moderate temporal variability. This configuration, summarized in Table 7, mimics the asynchronous and heterogeneous acquisition patterns encountered in operational weather stations and provides a more realistic testbed for evaluating forecasting models under non-uniform input conditions.
@@
Figure 5: Overview of the LNG Cargo Handling System (CHS).
Monash-SolarWind-practical (SolarWind)
@@ A.2 Implementation Details
All experiments were conducted using a single NVIDIA RTX 3090 GPU with 24GB memory. We used the Adam optimizer across all models, with initial learning rates per dataset for fair comparison: $10^{-4}$ for ETT1, ETT2, SolarWind and EPA, $10^{-2}$ for Weather, and $10^{-3}$ for the CHS dataset. Training was run for a maximum of 10 epochs with early stopping based on validation performance. For models with a stackable encoder block, the number of encoder layers was set to 2. For Transformer-based models, the model dimension $d_{\text{model}}$ was searched over $\{128,256,512\}$, and the feedforward expansion ratio $d_{\text{ff}}$ / $d_{\text{model}}$ was selected from $\{1,2,4\}$. We adopted most baseline implementations from the TimesNet (Wu et al., 2023) repository. CrossGNN (Huang et al., 2023), t-PatchGNN (Zhang et al., 2024), Hi-Patch (Luo et al., 2025), BiTGraph (Chen et al., 2024b), originally provided in a separate repository, were reimplemented and integrated into the TimesNet framework for consistency. The hyperparameters of these baselines were aligned with those specified in their original repositories. Since ContiFormer (Chen et al., 2023) is an ODE-based model with prohibitively long training times, we reimplemented it within the TimesNet framework by approximating its ODE solver, following the structure of TimeXer (Wang et al., 2024e), to ensure both efficiency and consistency. In our proposed ChannelTokenFormer (CTF), the number of channel tokens was selected from $\{1,2,3\}$, with the optimal value determined separately for each dataset.
@@
For regular time series baseline models, both training and testing were conducted on linearly interpolated data due to architectural constraints. To ensure a fair and consistent comparison, we preserved the original design of the baselines and trained them on interpolated inputs with MSE loss. In contrast, our proposed model, ChannelTokenFormer (CTF), is explicitly designed to operate on non-interpolated inputs that reflect the original, variable sampling periods of each channel. To ensure uniform input shape compatible with existing architectures, we apply forward-fill interpolation during dataset preprocessing; however, this step serves merely as a structural tool to meet input dimensionality requirements, not as a modeling assumption. During the channel-wise patch embedding phase, all forward-filled values are explicitly excluded: only the valid indices $L_{i}$, corresponding to truly observed data points, are used for computation. For irregular and missing handling time series baselines, we followed their original frameworks by providing an input mask, which was used for both data processing and MSE computation. This procedure is exactly consistent with the forward-fill interpolation–based approach. However, since ContiFormer does not conform to other irregular frameworks and was reimplemented for execution, it follows the same procedure as the regular baselines. To ensure fair evaluation, both the baselines and our model compute channel-aggregated metrics—CMSE and CMAE, as defined in Section 3—only over the valid indices $H_{i}$, which correspond to actually observed target values. Thus, although the input preprocessing methods differ, all models are evaluated on the same set of ground truth targets $(y_{1},\dots,y_{H_{i}})$. Note that under the practical setting, the number of valid test samples is inherently reduced, since the test set is subsampled according to the maximum sampling factor across channels. In summary, each model is trained under settings aligned with its architectural requirements, but all are evaluated uniformly based on their accuracy in predicting truly observed values.
@@
As described in Section 5, we set the input and prediction lengths for each dataset to reflect the temporal characteristics of the underlying signals. Instead of fixing the number of time steps across all channels, we define these lengths over a consistent time duration to accommodate channel-wise variations in both sampling period and dominant frequency components. For implementation and reporting convenience, however, we represent input and output lengths in terms of the number of steps corresponding to the channel with the smallest sampling period. The input and prediction lengths (in time steps, relative to the finest resolution channel) for each dataset are as follows: ETT1, ETT2: 192 (input) $\rightarrow$ [192, 336, 768, 1152] (prediction), SolarWind: 576 $\rightarrow$ [288, 576, 864, 1152], Weather: 144 $\rightarrow$ [144, 288, 576, 864], EPA: 96 $\rightarrow$ [96, 192, 288, 384], CHS: 120 $\rightarrow$ [120, 240, 480, 840].

### v4

@@ 5.1 Main Results
We evaluate forecasting performance using datasets that reflect realistic sampling heterogeneity across channels. Unlike standard baselines that rely on interpolation, our approach preserves the original sampling structure and operates directly on observed values. When we average CMSE and CMAE over all prediction lengths (Table 1), CTF delivers strong overall performance, achieving the best or second-best results across datasets. This underscores the benefit of avoiding interpolation when channels are temporally misaligned. In particular, CTF handles asynchrony more effectively than irregular modeling on most datasets. Detailed results for each prediction length and dataset are provided in Appendix E. As CTF uses frequency-domain information for dynamic patching, we also compare it against frequency-based approaches and still consistently achieve superior performance (see Appendix C.5 for more details).
@@
We further evaluate all models under a more challenging setting, where test-time inputs include block-wise missing intervals, beyond discrete missing values, explicitly indicated to the model, on top of channel-wise asynchronous sampling. This configuration encompasses all three core challenges of interest. In this case, all models receive inputs containing contiguous missing segments filled with zeros. But for a fair comparison, linear interpolation is applied to regular baselines to avoid underestimating their capability. Unlike other approaches, CTF can internally adjust input length by design, enabling the use of zero-patch masking to effectively handle missing regions. As shown in Table 2, our model consistently outperforms most baselines in average CMSE and CMAE for overall prediction lengths across different missing ratios, maintaining high performance even as the severity of missingness increases. This robustness stems from the use of random patch masking during training, which improves the model’s ability to generalize to incomplete patterns. Unlike other methods, CTF avoids distortion from artificially zero-filled values and preserves signal fidelity under block-wise missing conditions. Detailed results across datasets are provided in Appendix E. The corresponding details are provided in Appendix A.2.
@@ 7 Conclusion
A practical forecasting setting is introduced where three core real-world challenges occur simultaneously: channel-wise asynchronous sampling, block-wise missing intervals, and complex cross-channel dependencies. Instead of interpolation, our approach handles both missing segments and sampling gaps through masking and frequency-based dynamic patching. The unified mask-guided attention leverages channel tokens with read-only masking and cross-channel attention, consistently outperforming prior methods across real-world conditions. Our scalability analysis indicates that the masking-guided unified attention design remains computationally manageable for realistic channel counts and input lengths. However, ultra-high-dimensional regimes with thousands of channels (e.g., Traffic) will require additional mechanisms such as explicit channel sparsification or grouped channel representations, which we leave for future work. Beyond scaling, future work will also consider applying ChannelTokenFormer to broader domains, integrating multimodal signals, and improving interpretability of channel interactions, thereby expanding its utility in real-world forecasting.
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00345809, Research on AI Robustness Against Distribution Shift in Real-World Scenarios; and No. RS-2023-00222663, Center for Optimizing Hyperscale AI Models and Platforms). We sincerely thank HD Korea Shipbuilding & Offshore Engineering Co., Ltd. (HD KSOE), particularly the Decarbonization Ship Research Institute, for granting access to proprietary data under a confidentiality agreement and for supporting the secure data-sharing process that enabled this research.
@@ A.1 Dataset Configuration under Multi-Source Asynchrony
To further contextualize this setting, we note that widely used benchmarks already involve substantial preprocessing to obtain complete, regularly sampled grids. For example, the ETT datasets typically aggregate higher-frequency sensor streams for load-related channels into 15-minute or hourly averages. This procedure reduces the statistical complexity of the original processes and introduces its own form of synthetic regularity. Our practical variants do not create any new values; instead, they relax this enforced uniformity by selectively downsampling channels whose variability, assessed in the frequency domain, indicates sufficient oversampling. The resulting multi-rate sensing patterns more plausibly reflect real-world conditions while not materially altering the underlying temporal dynamics.
@@
Monash-SolarWind-practical (SolarWind)
@@
Weather-practical (Weather)
Weather-practical is constructed from the widely used Weather (Wu et al., 2021) benchmark. While the original dataset assumes uniform sampling periods across all channels, this assumption rarely holds in real-world environmental monitoring systems. In Weather-practical, we reorganize the channels to reflect heterogeneous sampling frequencies, based on the intrinsic temporal dynamics of each channel. For instance, fast-changing environmental channels such as wind velocity, precipitation, and solar radiation are sampled at 10-minute intervals, whereas more stable channels such as air pressure or vapor pressure are sampled at a coarser hourly rate. Mid-range channels such as air temperature and specific humidity are sampled at 20-minute or 30-minute intervals in accordance with their moderate temporal variability. This configuration, summarized in Table 7, mimics the asynchronous and heterogeneous acquisition patterns encountered in operational weather stations and provides a more realistic testbed for evaluating forecasting models under non-uniform input conditions.
@@ A.2 Implementation Details
All experiments were conducted using a single NVIDIA RTX 3090 GPU with 24GB memory. We used the Adam optimizer across all models, with initial learning rates per dataset for fair comparison: $10^{-4}$ for ETT1, ETT2, SolarWind and EPA, $10^{-2}$ for Weather, and $10^{-3}$ for the CHS dataset. Training was run for a maximum of 10 epochs with early stopping based on validation performance. For models with a stackable encoder block, the number of encoder layers was set to 2. For Transformer-based models, the model dimension $d_{\text{model}}$ was searched over $\{128,256,512\}$, and the feedforward expansion ratio $d_{\text{ff}}$ / $d_{\text{model}}$ was selected from $\{1,2,4\}$. Details of the hyperparameters used for each dataset are fully specified in the provided code repository  [^3] Code: https://github.com/jinkwan1115/ChannelTokenFormer.git. We adopted most baseline implementations from the TimesNet (Wu et al., 2023) repository. DUET (Qiu et al., 2025), TimeFilter (Hu et al., 2025), CrossGNN (Huang et al., 2023), t-PatchGNN (Zhang et al., 2024), Hi-Patch (Luo et al., 2025), BiTGraph (Chen et al., 2024b), originally provided in a separate repository, were reimplemented and integrated into the TimesNet framework for consistency. The hyperparameters of these baselines were aligned with those specified in their original repositories. In our proposed ChannelTokenFormer (CTF), the number of channel tokens was selected from $\{1,2,3\}$, with the optimal value determined separately for each dataset.
@@
For regular time series baseline models, both training and testing were conducted on linearly interpolated data due to architectural constraints. To ensure a fair and consistent comparison, we preserved the original design of the baselines and trained them on interpolated inputs with MSE loss. In contrast, our proposed model, ChannelTokenFormer (CTF), is explicitly designed to operate on non-interpolated inputs that reflect the original, variable sampling periods of each channel. To ensure uniform input shape compatible with existing architectures, we apply forward-fill interpolation during dataset preprocessing; however, this step serves merely as a structural tool to meet input dimensionality requirements, not as a modeling assumption. During the channel-wise patch embedding phase, all forward-filled values are explicitly excluded: only the valid indices $L_{i}$, corresponding to truly observed data points, are used for computation. For irregular and missing handling time series baselines, we followed their original frameworks by providing an input mask, which was used for both data processing and MSE computation. This procedure is exactly consistent with the forward-fill interpolation–based approach. To ensure fair evaluation, both the baselines and our model compute channel-aggregated metrics (CMSE and CMAE, as defined in Section 3) only over the valid indices $H_{i}$, which correspond to actually observed target values. Thus, although the input preprocessing methods differ, all models are evaluated on the same set of ground truth targets $(y_{1},\dots,y_{H_{i}})$. Note that under the practical setting, the number of valid test samples is inherently reduced, since the test set is subsampled according to the maximum sampling factor across channels. In summary, each model is trained under settings aligned with its architectural requirements, but all are evaluated uniformly based on their accuracy in predicting truly observed values. Some recent methods adopt a “drop-last” trick (Qiu et al., 2024) to improve performance. In our experiments, we do not apply this operation.
@@
As described in Section 5, we set the input and prediction lengths for each dataset to reflect the temporal characteristics of the underlying signals. Instead of fixing the number of time steps across all channels, we define these lengths over a consistent time duration to accommodate channel-wise variations in both sampling period and dominant frequency components. For implementation and reporting convenience, however, we represent input and output lengths in terms of the number of steps corresponding to the channel with the smallest sampling period. The input and prediction lengths (in time steps, relative to the finest resolution channel) for each dataset are as follows: ETT1, ETT2: 192 (input) $\rightarrow$ [192, 336, 768, 1152] (prediction), SolarWind: 576 $\rightarrow$ [288, 576, 864, 1152], Weather: 144 $\rightarrow$ [144, 288, 576, 864], EPA: 96 $\rightarrow$ [96, 192, 288, 384], CHS: 120 $\rightarrow$ [120, 240, 480, 840]. Additional details such as patch length for each channel in each public dataset are fully specified in the released code files.
@@
Algorithmic Details of Frequency-based Dynamic Patching and Tokenization
Algorithm 1 implements the frequency-based dynamic patching and tokenization procedure described below. Given that all channels are observed over the same time interval, we first compute the relative sampling period of each channel with respect to the smallest sampling period across channels. This determines how many samples (i.e., its effective sequence length) fall into this interval. For each channel, we then set the patch size primarily based on its dominant period. In practice, if the dominant period cannot be reliably estimated for a channel, one can instead define its patch size using only the relative sampling period, i.e., a sampling-aware patching rule. The resulting series for each channel is then partitioned into patches, and we discard patches that are either entirely unobserved during test time or randomly skipped with probability $q$ during training, so that the token sequence reflects only informative segments. Each remaining patch is linearly projected and enriched with positional embeddings, after which a small set of learnable channel tokens is appended and a shared channel embedding is added to all tokens from the same channel. This yields a unified token sequence that encodes heterogeneous sampling periods and channel-wise frequency characteristics without relying on interpolation.
Algorithm 1 Frequency-based Dynamic Patching and Tokenization
1:Multivariate series $X\in\mathbb{R}^{N\times L}$; sampling periods $\{s_{i}\}_{i=1}^{N}$; dominant periods $\{D_{i}\}_{i=1}^{N}$; horizon $H$; mask prob. $q$; #channel tokens $m$
2:Token sequence $S$
[…]
5:for $i=1$ to NN do
6:  $s^{\prime}_{i}\leftarrow s_{i}/s_{\min}$ $\triangleright$ relative sampling period
7:  $L_{i}\leftarrow L/s^{\prime}_{i}$ $\triangleright$ valid length of channel $i$
8:  $p_{i}\leftarrow D_{i}/s^{\prime}_{i}$ $\triangleright$ patch length in samples
9:  $K_{i}\leftarrow L_{i}/p_{i}=L/D_{i}$ $\triangleright$ number of patches
10:  Partition $X^{(i)}\in\mathbb{R}^{L_{i}}$ into $\{P^{(i)}_{j}\}_{j=1}^{K_{i}}$, where $P^{(i)}_{j}\in\mathbb{R}^{p_{i}}$
11:  $T^{(i)}\leftarrow[\,]$ $\triangleright$ local tokens kept after masking
12:  for $j=1$ to $K_{i}$ do
13:   if $P^{(i)}_{j}$ is fully unobserved (test time) or $\mathrm{Bernoulli}(q)=1$ (training) then $\triangleright$ patch masking
14:     continue
15:   end if
16:   $T^{(i)}_{j}\leftarrow W_{p_{i}}P^{(i)}_{j}+e^{\mathrm{pos}}_{i,j}$ $\triangleright$ $W_{p_{i}}\!\in\!\mathbb{R}^{d\times p_{i}},\,T^{(i)}_{j}\!\in\!\mathbb{R}^{d}$
17:   Append $T^{(i)}_{j}$ to $T^{(i)}$
18:  end for
19:  Initialize learnable channel tokens $\{C^{(i)}_{k}\}_{k=1}^{m}$, $C^{(i)}_{k}\in\mathbb{R}^{d}$
20:  $S^{(i)}\leftarrow\big[T^{(i)},\,C^{(i)}_{1},\dots,C^{(i)}_{m}\big]+e^{\mathrm{chn}}_{i}$ $\triangleright$ $e^{\mathrm{chn}}_{i}$ is broadcast to all tokens
21:  Append $S^{(i)}$ to $S$
22:end for
23:return $S$
@@
For controlled missingness injection, we vary the overall missing ratio and apply block-wise masking by independently sampling contiguous blocks for each channel and each sample. This procedure mimics naturally occurring block-wise gaps caused by sensor malfunctions or maintenance events in real deployments. It produces a range of challenging, randomly located missing blocks that are applied identically to all compared methods at test time.